OpenStudy (anonymous):

(3,8);x+7y=8 Y=mx+b I need to gigure out Y=

OpenStudy (anonymous):

What's this x+7y=8 ?

OpenStudy (anonymous):

its equation and parellel and I dont get these problems

OpenStudy (anonymous):

7y = -x + 8 y = -1/7 x + 8/7 slope is -1/7 y = -1/7 x + b 8 = -3/7 + b 59/7 = b equation is y = -1/7 x + 59/7

OpenStudy (anonymous):

wow .... I get cofused once it goes to a fraction

OpenStudy (anonymous):

May be you should do some fraction problems on Khan Academy site. To get you in the groove.

OpenStudy (anonymous):

@karate it should be a negative slope

OpenStudy (karatechopper):

o srry yes it should my bad i am multitasking

OpenStudy (anonymous):

You may want to delete your answer. So as to keep it clean.

OpenStudy (karatechopper):

k